Transverse waves on a string obey the equation -where u(x,t) is the time Cu (a) Ifat t=0, the transverse displacement u(x,0) =_and__ = 0 find u(x,t)Use the 1+x Cx general solution of the wave equation. (b) Ifthe two ends ofthe string are fixed at x=0 and x=a and at t=0 u(x,0)-0, adx,t= 0) = dx(a-x) , use separation of variables to obtain u(x,t). Cx
Answer #1

We have the differential equation for the transverse wave is

d2u/dx2=(1/v2)d2u/dt2. d is the partial differential operator

We can write above v equation as

X"t=(1/v2)(t"x).

x"/x=t"/t=v2.

Finally two equations become

X"-v2X=0. And t"-v2t=0.

And solutions become from equations D2-v2=0,roots D=+v,-v. And general solution become

u(x,t)=(c1evx+c2e-vx)(c3evt+c4e-vt). Transverse displacement.

Where x is the position,v is the wave speed and t is the time.

(a). Given initial conditions are. u(x,0)=(1/(1+x2)). And du/dx=0.

u(x,0)=(c1evx+c2e-vx)(c3+c4)=(1/(1+x2))-->c3+c4=(1/(1+x2))/(2c1evx)

du/dx=(c1vevx-c2ve-vx)(c3evt+c4e-vt)=0.

That means

V(c1evx-c2e-vx)=0

C1evx=c2e-vx.

C2=c1e2vx.

Finally general solution become if we are taking coefficients as zeros

u(x,t)=2evx(1/(1+x2)/(2exv)=(1/(1+x2)).

(b).Now,given initial conditions are u(x,0)=0. And substitute in the general solution

u(x,0)=c3+c4=0 an c3=-c4.

u(0,t), u(a,t) are fixed.means u(0,t)=0 and u(a,t)=0. Fixed means displacement will become zero.

u(0,t)=(c1+c2)=0 and c1=-c2.

And u(a,t)=c1eax+c2e-ax=0.c2=-c1e2ax.

du(x,t=0)/dx=Ax(a-x)=(c1evx-c2e-vx)(c3+c4).

So,Final soltiuon by seperating x and t variables is

u(x,t)=-e2ax(Ax(a-x))=-e2at(At(a-t))